# GastCoCo: Graph Storage and Coroutine-Based Prefetch Co-Design for Dynamic Graph Processing

### v3

@@ Abstract.
An efficient data structure is fundamental to meeting the growing demands in dynamic graph processing. However, the dual requirements for graph computation efficiency (with contiguous structures) and graph update efficiency (with linked list-like structures) present a conflict in the design principles of graph structures. After experimental studies of existing state-of-the-art dynamic graph structures, we observe that the overhead of cache misses accounts for a major portion of the graph computation time. This paper presents GastCoCo, a system with graph storage and coroutine-based prefetch co-design. By employing software prefetching via stackless coroutines and introducing a prefetch-friendly data structure CBList, GastCoCo significantly alleviates the performance degradation caused by cache misses. Our results show that GastCoCo outperforms state-of-the-art graph storage systems by 1.3$\times$ - 180$\times$ in graph updates and 1.4$\times$ - 41.1$\times$ in graph computation.
@@
The source code, data, and/or other artifacts have been made available at https://github.com/GorgeouszzZ/GastCoCo.
@@ 1. Introduction
(b) CPU cache stall proportion.
@@
Figure 1. The execution time for graph algorithms (and graph updates) and the CPU cache stall proportion on different data structures. T.O. means that the graph updates cannot be finished in a reasonable time (much longer than 48 hours).
@@
With the increase in data scale and diversification of data types, dynamic graph processing has become a critically important issue in various domains such as e-commerce, financial technology, and social networks (Azadifar et al., 2022; Tu et al., 2023; Zamini et al., 2022). Specifically, dynamic graph processing involves executing graph computation algorithms on graphs that experience frequent structural changes, such as millions of edge updates per hour (Sahu et al., 2017; He et al., 2023). For instance, there are more than $40$ million behaviors per day (Ali, 2023) between users and items on Taobao (Tao, 2024). The platform should employ a combination of graph analytics and interactive graph traversals for fraud detection or interactive pattern mining (Ding et al., 2020). Such applications on dynamic graphs require not only efficient graph computations but also supporting the large volume of updates per second (Gupta et al., 2014; Qiu et al., 2018; Sharma et al., 2016).
@@
Regrettably, the need for computation efficiency and frequent update support presents a contradictory challenge, leading to a dilemma in designing storage structures for dynamic graphs. Take the widely used graph storage structures, Adjacency List (AL) and Compressed Sparse Row (CSR)  (Saad, 2003), as an example. AL and CSR are two extremes in graph storage: CSR is completely contiguous in memory, while AL is highly discontiguous in memory. AL is a classic graph data structure where the neighbors of each vertex are stored in a linked list. Since the nodes in the linked list are non-contiguous memory fragments connected by pointers, it is easy to insert/delete nodes without data movements on the AL structure, which is very suitable for graph updates. On the other hand, CSR has been widely used for graph storage in static graphs because it is both memory efficient and graph computation efficient. Figure1a reports the execution time for graph computations (PageRank, SSSP, Triangle Count, and Edge Query) and updates (Insertion) under single thread on the LiveJournal social network dataset (Backstrom et al., 2006). We could observe that CSR is more efficient than AL on graph algorithms, but spends more time for updating graphs (Graph update experiment on CSR cannot finish in 48 hours).
@@
To effectively support both graph computation and graph updates, some efforts propose designing graph structures that maintain a moderate degree of contiguity (Zhu et al., 2020; Fuchs et al., 2022; Pandey et al., 2021; Ediger et al., 2012; Kumar and Huang, 2020). For example, Adjacency Array (AA)  (Zhu et al., 2020) and Block-based Adjacency List (BAL)  (Ediger et al., 2012) are such compromised storage structures, where AA stores all neighbors of each vertex in a separate array and BAL stores the neighbors of each vertex in a block-based linked list (i.e., a node contains more than one neighbors). The degrees of contiguity in these two structures fall between CSR and AL. More recently, more complex and detailed implementations, e.g., the state-of-the-art dynamic graph storage systems like GraphOne (Kumar and Huang, 2020), Stinger (Ediger et al., 2012), Terrace (Pandey et al., 2021), and Sortledton (Fuchs et al., 2022), maintain the graphs with a linked data structure fragmented in memory with various degrees of contiguity (more akin to CSR or AL) as a halfway house. However, such a halfway-house solution can only balance the performance of graph computation and graph updates which sacrifices the efficiency of graph computation to enhance graph update throughput, or vice versa.
@@
To avoid differences in optimization and implementation across systems, we use uniformly implemented simple data structures for the cache stall proportion experiments. As representatives of ”a halfway-house solution,” BAL and AL exhibit a significant proportion ($71.38\%$ to $96.21\%$) of cache miss overhead in graph computation tasks.This motivates us to seek a novel perspective to improve both the efficiency and dynamic graph support of dynamic graph processing: Can we mitigate the overhead of cache misses from the linked data structures to improve the efficiency for both computation and updating?
@@
To answer the question, we consider a co-design of the prefetching techniques and graph structures. By prefetching techniques, we can load data that will be accessed later from memory into the CPU cache to reduce the data access latency. Prefetching techniques encompass both hardware prefetching, where the CPU automatically loads data and instructions, and software prefetching, which involves programmers explicitly inserting instructions into the code to prefetch data. Although hardware prefetching techniques are commonly equipped on standard hardware, software prefetching techniques have recently demonstrated their effectiveness in many applications  (Callahan et al., 1991; He et al., 2020; Zhi et al., 2024). When it comes to the practical dynamic graph processing domain, there are still challenges.
@@
Our approach. This paper presents GastCoCo, an in-memory system with GrAph STorage and COroutine-based prefetch CO-design, which is designed for dynamic graph processing applications to alleviate the performance overhead caused by cache misses. ①In terms of hardware prefetching, the prefetching effectiveness varies across different data structures. GastCoCo proposes a prefetch-aware data structure CBList for dynamic graph processing. ②Software prefetching is used to compensate for cases where hardware prefetching fails, so the benefits of applying software prefetching are our focus. GastCoCo develops a set of hybrid prefetching strategies to avoid software prefetching and hardware prefetching from fetching reduplicated data. To minimize the overhead of applying software prefetching and to increase its prefetch success rate, GastCoCo leverages C++20 (ISO/IEC, 2017) stackless coroutines to prefetch the graph data that the program requires to access, benefiting both graph computation and graph updates by mitigating the cache miss overhead. ③Finally, GastCoCo designs a set of adaptive coroutine scheduling and task allocation strategies tailored to different graph tasks, allowing graph processing using software prefetching to perform better on GastCoCo.
@@ 2.1. Graph Operations and Data Access Patterns
Processing edges can be considered as first locating the edges via the source vertex $v_{src}$ and then processing the neighbors of $v_{src}$. Since locating the edges via the source vertex can be achieved by read_vertex(v), our focus is exclusively on the process of neighbors.
@@
Sequential Data Access. Some tasks require sequential data access, i.e., to access the storage structure in a fixed order (e.g., the lexicographical order). If the structure is physically contiguous in memory, sequential data access can be regarded as sequential memory access. For example, scan_vertices() and scan_edges($v_{src}$) can be regarded as sequential data access or as sequential memory access on arrays.
Random Data Access. Random data access refers to accessing the storage structure (or part) in a random order. Since the access has been unpredicted, performing random data access on either contiguous or non-contiguous memory can be considered as random memory access. Correspondingly, scan_vertices(cond) can be regarded as random data access due to its unpredictable condition $cond$. read_vertex(v) and read_edge($v_{src}$,$v_{dst}$) also access data in a random order.
@@ 2.2. Hardware Prefetching in Graph Processing
Hardware prefetching is a predefined mechanism in the processor that watches the stream of data being requested by the executing program, recognizes the next few elements that the program might need, and prefetches the data into the processor’s cache beforehand (Baer and Chen, 1991). However, its effectiveness is affected by different data structures. In general, to support efficient graph computation and updates, dynamic graph storage often incorporates both contiguous and non-contiguous memory structures. Specifically, Array-like structures (Saad, 2003; Wheatman and Xu, 2018; Leo and Boncz, 2021) use contiguous memory, while Linked List-like (Fuchs et al., 2022; Kumar and Huang, 2020; Ediger et al., 2012) and Tree-like structures (Dhulipala et al., 2019; Pandey et al., 2021; Fuchs et al., 2022) use non-contiguous memory.
@@
Hardware prefetching in Array-like structures. The data in an Array-like structure is stored contiguously in the main memory. As shown in Figure 2 1, when scan_edges($v_{src}$) is performed on an array, the hardware prefetchers prefetch the next one or several segments adjacent to the requested segment into the cache from the main memory. Here a segment refers to a cache line (segment size matches cache line size) (Smith, 1978; Baer and Chen, 1991; Falsafi and Wenisch, 2014). Therefore, when performing sequential memory access on array-like structures (e.g., scan_edges($v_{src}$) in CSR), most of the required data will be hit in the CPU cache. However, if the operation is randomly accessed, most of the data will be missed in cache.
@@
Hardware prefetching in Linked List-like structures. The data stored in the Linked List-like structure is discontiguous, and the data is connected by pointers in memory. Under the sequential data access pattern, the data to be accessed next can be known in advance via the fixed access paths. Modern hardware prefetchers will prefetch data in the sequential data access pattern by jump-pointer mechanism (Luk and Mowry, 1996; Collins et al., 2002; Roth et al., 1998; Roth and Sohi, 1999; Falsafi and Wenisch, 2014). In Figure 2 2, the jump-pointer mechanism prefetches Node $2$ when the processor reads Node $1$, preventing a cache miss when the processor reads Node $2$. However, such a jump-pointer mechanism does not work for the head of the linked list. For instance, upon initially entering the linked list in Figure 2 2, Node $1$ can not be prefetched in time, leading to a cache miss. Regarding the occurrence of cache misses, we provide the following explanation: i) The first node is not located in the caches with a high possibility while initially accessing a linked list. Therefore, it results in a cache miss (in fact, the front few nodes might trigger cache misses, depending on how many nodes the jump-pointer mechanism prefetches ahead). ii) Even if the memory address of the next node is far from the current block, the jump-pointer mechanism can load the afterward nodes in advance while traversing on a linked list as shown in Figure 2. Therefore, fewer cache misses occur in the subsequent process.
@@
Hardware prefetching in Tree-like structures. Several previous works, like Aspen (Dhulipala et al., 2019), and Terrace (Pandey et al., 2021), utilize tree-like structures to store the graphs for high-degree vertices. In these structures, there is little inherent spatial locality between accessed nodes since they are dynamically allocated from the heap and can have arbitrary addresses (Luk and Mowry, 1996). The hardware prefetching will no longer work for such structures (He et al., 2020; Falsafi and Wenisch, 2014) because blocks may have more than one pointer. That is, hardware prefetchers cannot predict which path you are about to take next. As a result, both scan_edges($v_{src}$) and read_edge($v_{src}$,$v_{dst}$) (sequential and random data access in a tree) result in a large number of cache misses.
@@
In summary, hardware prefetching mechanisms provide more direction for our data structure design. However, the unexpected performance of hardware prefetching inspires us to employ other modern techniques to reduce the time of cache misses or even eliminate some of the cache misses from data fetching, such that a fragmented data structure for dynamic graph computing could also achieve better performance.
@@ 2.3. Software Prefetching via Coroutines
Although hardware prefetching could relieve the access time for contiguous data, it performs badly for structural data like graphs. In response, recent efforts in the database community utilize coroutines (de Moura and Ierusalimschy, 2009; Jonathan et al., 2018) to mitigate cache misses. For example, prior works (Jonathan et al., 2018; Chen et al., 2007; Psaropoulos et al., 2017; He et al., 2020) propose to optimize the overhead of cache misses caused by random memory access. This solution is implemented by software prefetching via coroutines. Software prefetching techniques (Jonathan et al., 2018; Chen et al., 2007; Psaropoulos et al., 2017; Koçberber et al., 2015) leverage workload semantics to issue prefetch instructions (Intel Corporation, 2016) to explicitly bring data into CPU caches (He et al., 2020). Unlike hardware prefetching, software prefetching is controlled by the instructions written by programmers explicitly.
This solution enables the overlapping of data loading with computation. When a task requires data, the data can be explicitly loaded using prefetch instructions. However, in consideration of the delay as the data moves from memory to cache, CPU switches to another task to compute to overlap data loading time. In such a solution, we need a tool to finish the switching between tasks. A naive approach would be to use multithreading for switching, but the overhead of thread switching far exceeds the cost of LLC cache miss. To this end, we need new techniques to switch between tasks.
C++20 Stackless Coroutine are special functions that can be suspended voluntarily and resumed later at a low cost during their execution (de Moura and Ierusalimschy, 2009; Conway, 1963; He et al., 2020). The efficient suspension and resumption of stackless coroutines make it possible to implement the interleaving execution mode. To be more specific, when issuing a software prefetching instruction, we can suspend the coroutine, switch to another coroutine for computing, and later resume the previous coroutine to continue the corresponding task. Stackless coroutines are essentially state machines, i.e., coroutine switching simply involves changing the instruction pointer register. All coroutines within the same physical thread share the system stack, and there is no need to save the values of registers explicitly. So the stackless coroutines in the C++20 (ISO/IEC, 2017) standard exhibit low overhead in terms of construction and context switching (cheaper than a last-level cache miss) (He et al., 2020), and GastCoCo is constructed based on C++20 stackless coroutines.
Hardware prefetching is an automatic prefetching technique executed by the processors, but in the scenarios of dynamic graph processing, in many cases (as discussed in subsection 2.2) hardware prefetching does not work. We apply software prefetching to complement hardware prefetching, achieving a more effective prefetching result. But using software prefetching via coroutines also incurs certain overheads. Therefore, how to implement effective software prefetching in dynamic graph domain is the focus of our work.

### v4

@@ Abstract.
An efficient data structure is fundamental to meeting the growing demands in dynamic graph processing. However, the dual requirements for graph computation efficiency (with contiguous structures) and graph update efficiency (with linked list-like structures) present a conflict in the design principles of graph structures. After experimental studies of state-of-the-art dynamic graph structures, we observe that the overhead of cache misses accounts for a major portion of the graph computation time. This paper presents GastCoCo, a system with graph storage and coroutine-based prefetch co-design. By employing software prefetching via stackless coroutines and designing a prefetch-friendly data structure CBList, GastCoCo significantly alleviates the performance degradation caused by cache misses. Our results show that GastCoCo outperforms state-of-the-art graph storage systems by 1.3$\times$ - 180$\times$ in graph updates and 1.4$\times$ - 41.1$\times$ in graph computation.
PVLDB Reference Format:
Hongfu Li, Qian Tao, Song Yu, Shufeng Gong, Yanfeng Zhang, Feng Yao, Wenyuan Yu, Ge Yu, and Jingren Zhou. PVLDB, 17(13): XXX-XXX, 2024.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 17, No. 13 ISSN 2150-8097.
doi:XX.XX/XXX.XX
@@
The source code, data, and/or other artifacts have been made available at https://github.com/GorgeouszzZ/GastCoCo. 11footnotetext: Hongfu Li and Qian Tao contributed equally.
@@ 1. Introduction
(b) CPU cache stall count.
@@
Figure 1. The execution time for graph algorithms (and graph updates) and the CPU cache stall count on different data structures (T.O.: graph updates cannot finish in 24 hours).
@@
With the increase in data scale and diversification of data types, dynamic graph processing has become a critically important issue in various domains such as e-commerce, financial technology, and social networks (Azadifar et al., 2022; Tu et al., 2023; Zamini et al., 2022). Specifically, dynamic graph processing involves executing graph computation algorithms on graphs that experience frequent structural changes, such as millions of edge updates per hour (Sahu et al., 2017; He et al., 2023). For instance, there are more than $400$ million behaviors per day (Ali, 2023) between users and items on Taobao (Tao, 2024). The platform should employ a combination of graph analytics and interactive graph traversals for fraud detection or interactive pattern mining (Ding et al., 2020). Such applications on dynamic graphs require not only efficient graph computations but also supporting the large volume of updates per second (Gupta et al., 2014; Qiu et al., 2018; Sharma et al., 2016).
@@
Regrettably, the need for computation efficiency and frequent update support presents a contradictory challenge, leading to a dilemma in designing storage structures for dynamic graphs. Take the widely used graph storage structures, Adjacency List (AL) and Compressed Sparse Row (CSR)  (Saad, 2003), as an example. AL is a graph data structure where the neighbors of each vertex are stored in a linked list. Since the nodes in the linked list are non-contiguous memory fragments connected by pointers, it is easy to insert/delete nodes without data movements on the AL structure, which is suitable for graph updates. On the other hand, CSR has been widely used for graph storage in static graphs due to its memory and computational efficiency. Figure1a reports the execution time for graph computations (PageRank, SSSP, Triangle Count, and Edge Query) and updates (Insertion) under a single thread on the LiveJournal dataset (Backstrom et al., 2006). We could observe that CSR is more efficient than AL on graph algorithms, but spends more time for updating graphs.
@@
To effectively support both graph computation and graph updates, some efforts propose designing graph structures that maintain a moderate degree of contiguity (Zhu et al., 2020; Fuchs et al., 2022; Pandey et al., 2021; Ediger et al., 2012; Kumar and Huang, 2020). For example, Adjacency Array (AA) (Zhu et al., 2020) stores all neighbors of each vertex in a separate array and Block-based Adjacency List (BAL) (Ediger et al., 2012) stores the neighbors of each vertex in a block-based linked list (i.e., a node contains more than one neighbors). The degrees of contiguity in these two structures fall between CSR and AL. More recently, more complex and detailed implementations, e.g., the state-of-the-art dynamic graph storage systems like GraphOne (Kumar and Huang, 2020), Stinger (Ediger et al., 2012), Terrace (Pandey et al., 2021), and Sortledton (Fuchs et al., 2022), maintain the graphs with a linked data structure fragmented in memory with various degrees of contiguity as a halfway house. However, such a halfway-house solution can only balance the performance of graph computation and graph updates, which sacrifices the efficiency of graph computation to enhance graph update throughput, or vice versa.
@@
To avoid differences in optimization and implementation across systems, we use uniformly implemented simple data structures for the cache stall count experiments in Figure 1b. As representatives of “halfway-house solutions”, BAL and AL exhibit a large number of cache misses in graph computation tasks. This motivates us to seek a novel perspective to improve both the efficiency and dynamic graph support of dynamic graph processing: Can we mitigate the overhead of cache misses from the linked data structures to improve the efficiency for both computation and updating?
@@
To answer the question, we consider a co-design of the prefetching techniques and graph structures. Prefetching techniques allow us to load data, which will be accessed later, from memory into the CPU cache, thereby reducing the data access latency. These techniques consist of both hardware prefetching, where the CPU automatically loads data and instructions, and software prefetching, where programmers explicitly insert instructions into the code to prefetch data. Although hardware prefetching techniques are commonly equipped on standard hardware, software prefetching techniques have recently demonstrated their effectiveness in many applications  (Callahan et al., 1991; He et al., 2020; Zhi et al., 2024). However, when it comes to the practical dynamic graph processing domain, there are still challenges.
@@
Our approach. This paper presents GastCoCo, an in-memory system with GrAph STorage and COroutine-based prefetch CO-design, which is designed for dynamic graph processing applications to alleviate the performance overhead caused by cache misses. ①In terms of hardware prefetching, the prefetching effectiveness varies across different data structures. GastCoCo proposes a prefetch-aware data structure CBList for dynamic graph processing. ②Software prefetching is used to compensate for cases where hardware prefetching fails, so the benefits of applying software prefetching are our focus. GastCoCo develops a set of hybrid prefetching strategies to avoid fetching reduplicated data by both software and hardware prefetching. To minimize the overhead of applying software prefetching and increase the prefetch success rate, GastCoCo leverages C++20 (ISO/IEC, 2017) stackless coroutines to prefetch the graph data that the program requires to access, benefiting both graph computation and graph updates by mitigating the cache miss overhead. ③Finally, GastCoCo designs a set of adaptive coroutine scheduling and task allocation strategies tailored to different graph tasks, allowing graph processing using software prefetching to perform better on GastCoCo.
@@ 2.1. Graph Operations and Data Access Patterns
Processing edges can be considered as first locating the edges via the source vertex $v_{src}$ and then processing the neighbors of $v_{src}$. Since locating the source vertex can be achieved by read_vertex(v), our focus is exclusively on the process of neighbors.
@@
Sequential Data Access. Some tasks require sequential data access, i.e., to access the storage structure in a fixed order (e.g., the lexicographical order). If the structure is physically contiguous in memory, sequential data access can be regarded as sequential memory access. For example, scan_vertices() and scan_edges($v_{src}$) can be regarded as sequential data access or memory access on arrays.
Random Data Access. Random data access refers to accessing the storage structure (or part) in a random order. Since the access is unpredicted, performing random data access on either contiguous or non-contiguous memory can be considered as random memory access. Correspondingly, scan_vertices(cond) can be regarded as random data access because $cond$ is unpredictable. read_vertex(v) and read_edge($v_{src}$,$v_{dst}$) also access data in a random order.
@@ 2.2. Hardware Prefetching in Graph Processing
Hardware prefetching is a predefined mechanism in the processor that, based on the data stream requested by the running program, identifies and prefetches the subsequent elements that the program might need, loading them into the processor’s cache beforehand (Baer and Chen, 1991). However, its effectiveness is affected by different data structures. Dynamic graph storage generally incorporates contiguous and non-contiguous memory structures to support efficient graph computation and updates. Specifically, Array-like structures (Saad, 2003; Wheatman and Xu, 2018; Leo and Boncz, 2021) use contiguous memory, while Linked List-like (Fuchs et al., 2022; Kumar and Huang, 2020; Ediger et al., 2012) and Tree-like structures (Dhulipala et al., 2019; Pandey et al., 2021; Fuchs et al., 2022) use non-contiguous memory.
@@
Hardware prefetching in Array-like structures. The data in an Array-like structure is stored contiguously in the main memory. As shown in Figure 2 1, when scan_edges($v_{src}$) is performed on an array, the hardware prefetchers prefetch the next several segments adjacent to the requested segment into the cache from the main memory. Here, the size of a segment matches that of a cache line  (Smith, 1978; Baer and Chen, 1991; Falsafi and Wenisch, 2014). Therefore, when performing sequential memory access on array-like structures (e.g., scan_edges($v_{src}$) in CSR), most of the required data will be hit in the CPU cache. However, if the operation randomly accesses data, most of the data will be missed.
@@
Hardware prefetching in Linked List-like structures. The data stored in the Linked List-like structure is discontiguous, and pointers in memory connect the data. Under the sequential data access pattern, the data to be accessed next can be known in advance via the fixed access paths. Modern hardware prefetchers prefetch data in the sequential data access pattern by jump-pointer mechanism (Luk and Mowry, 1996; Collins et al., 2002; Roth et al., 1998; Roth and Sohi, 1999; Falsafi and Wenisch, 2014). In Figure 2 2, the jump-pointer mechanism prefetches Node $2$ when the processor reads Node $1$, preventing a cache miss when the processor reads Node $2$. However, such a jump-pointer mechanism does not work for the head of a linked list. For instance, upon initially entering the linked list in Figure 2 2, Node $1$ can not be prefetched, leading to a cache miss. We offer the following explanation regarding the occurrence of cache misses: i) The first node is not located in the caches with a high possibility while accessing a linked list. Therefore, it results in a cache miss (in fact, the front few nodes might trigger cache misses, depending on how many nodes the jump-pointer mechanism prefetches ahead). ii) Even if the memory address of the next node is far from the current block, the jump-pointer mechanism can load the afterward nodes in advance while traversing on a linked list as shown in Figure 2. Therefore, fewer cache misses occur in the subsequent process.
@@
Hardware prefetching in Tree-like structures. Several previous works, like Aspen (Dhulipala et al., 2019) and Terrace (Pandey et al., 2021), utilize tree-like structures to store the high-degree vertices. In these structures, there is little inherent spatial locality between accessed nodes since they are dynamically allocated from the heap and can have arbitrary addresses (Luk and Mowry, 1996). The hardware prefetching no longer works for such structures (He et al., 2020; Falsafi and Wenisch, 2014) as blocks may contain more than one pointer, and hardware prefetchers cannot predict the path that will be selected. Therefore, both scan_edges($v_{src}$) and read_edge($v_{src}$,$v_{dst}$) (sequential and random data access in a tree) result in a large number of cache misses.
@@
In summary, hardware prefetching mechanisms provide more direction for our data structure design. However, the unexpected performance of hardware prefetching inspires us to employ other modern techniques to eliminate some of the cache misses from data fetching, such that a fragmented data structure for dynamic graph computing could also achieve better performance.
@@ 2.3. Software Prefetching via Coroutines
Although hardware prefetching relieves the access time for contiguous data, it performs badly on graphs. In response, recent efforts utilize coroutines (de Moura and Ierusalimschy, 2009; Jonathan et al., 2018) to mitigate cache misses. For example, prior works (Jonathan et al., 2018; Chen et al., 2007; Psaropoulos et al., 2017; He et al., 2020) propose optimizing the overhead of cache misses caused by random memory access, by introducing software prefetching via coroutines. Software prefetching techniques (Jonathan et al., 2018; Chen et al., 2007; Psaropoulos et al., 2017; Koçberber et al., 2015) leverage workload semantics to issue prefetch instructions (Intel Corporation, 2016) to explicitly bring data into CPU caches (He et al., 2020).
This solution enables the overlapping of data loading and computation. Data can be explicitly loaded using prefetch instructions as needed. However, given the delay involved in moving data from memory to cache, the CPU switches to another task to compute to overlap with data loading time. We need a tool to finish the switching between tasks. While the multithreading approach is straightforward, it has much heavy overhead on thread switching relative to the cost of LLC cache miss. To this end, we need new techniques to switch between tasks.
C++20 Stackless Coroutine are special functions that can be suspended voluntarily and resumed later at low cost during the execution (de Moura and Ierusalimschy, 2009; Conway, 1963; He et al., 2020). The efficient suspension and resumption of stackless coroutines make it possible to implement the interleaving execution mode. Specifically, when issuing a software prefetching instruction, we can suspend the coroutine, switch to another coroutine for computing, and later resume the previous coroutine to continue the corresponding task. Stackless coroutines are essentially state machines, i.e., coroutine switching simply involves changing the instruction pointer register. All coroutines within a physical thread share the system stack, and it is unnecessary to save the states of registers explicitly. The stackless coroutines in C++20 (ISO/IEC, 2017) standard exhibit low overhead in terms of construction and context switching (He et al., 2020), and GastCoCo is constructed based on C++20 stackless coroutines.
Hardware prefetching is an automatic prefetching technique executed by the processors but shows inefficiency in many cases (as discussed in subsection 2.2). We apply software prefetching to complement hardware prefetching, achieving a more effective performance. Switching between coroutines can cause considerable cost, and our work focuses on how to implement effective software prefetching for dynamic graphs.